In a titration, a student found that 42.6 mL of a KOH solution were required to neurtralize 28.9 mL of a 0.160 M sulfuric acid solution. Determine the molarity of the KOH solution. SHOW WORK.

2KOH(aq) + H2SO4(aq) ------> K2SO4(aq) + 2H2O(l)
Answer #1

Balanced chemical equation is:

1 H2SO4 + 2 KOH ---> 1 K2SO4 + 2 H2O

According to balanced reaction:

2*number of mol of H2SO4 =1*number of mol of KOH

2*M(H2SO4)*V(H2SO4) =1*M(KOH)*V(KOH)

2*0.16*28.9 = 1*M(KOH)*42.6

M(KOH) = 0.217M

Answer: 0.217 M